# WATonoBus: An All Weather Autonomous Shuttle

### v1

@@ Abstract
Autonomous vehicle all-weather operation poses significant challenges, encompassing modules from perception and decision-making to path planning and control. The complexity arises from the need to address adverse weather conditions like rain, snow, and fog across the autonomy stack. Conventional model-based and single-module approaches often lack holistic integration with upstream or downstream tasks. We tackle this problem by proposing a multi-module and modular system architecture with considerations for adverse weather across the perception level, through features such as snow covered curb detection, to decision-making and safety monitoring. Through daily weekday service on the WATonoBus platform for almost a year, we demonstrate that our proposed approach is capable of addressing adverse weather conditions and provide valuable learning from edge cases observed during operation.
Index Terms:
Autonomous shuttle, dependable perception, dependable localization, adverse weather, pullover and merging, system architecture.
@@ I Introduction
Autonomous shuttle bus technology has been steadily advancing in recent years [1]. The introduction of self-driving shuttle buses that can operate along predefined routes without the need for human drivers will enable a new generation of the first- and last-mile transportation solutions. Autonomous shuttle buses have the potential to reduce traffic congestion and improve traffic efficiency. By eliminating the need for drivers to operate the vehicle, human errors are minimized and road safety will therefore increase due to the corresponding reduction in collision frequency [2]. Through collaborations among universities, government bodies, and industry partners, numerous cities have initiated pilot programs to test and deploy autonomous shuttle buses in controlled environments [3, 4, 5]. Of these, most are simply exploratory and are motivated by short-term objectives that include, for example, the collection of autonomous driving data, the evaluation of economic benefits, and the opportunity to increase public exposure to autonomous driving technology. Although autonomous shuttle buses provide many benefits in comparison to traditional transportation methods, the task of matching or exceeding the performance of a human driver under dynamic weather and lighting conditions remains a challenge. One related safety standard has been drafted to provide qualitative guidelines for testing the performance of autonomous shuttle bus projects [6], however, the multi-seasonal autonomous shuttle bus projects are still rarely found, posing challenges on the practicability of this technology.
@@ I-A Related works
In December 1997, Schiphol Airport in Netherland unveiled one of the early instances of an ”automated people mover” system, known as the ParkShuttle [7]. This pivotal development laid foundation for testing autonomous driving capabilities within a public environment. Currently, numerous global companies have successfully manufactured pilot self-driving shuttle buses for commercial purposes. These firms are Apollo, EasyMile, Navya, and Olli [8, 9, 10]. WEpod project was initiated in the Netherlands with the aim of enhancing our understanding of autonomous vehicles (AVs) [11]. Two driverless shuttles navigated through the Wageningen University campus, smoothly interacting with regular traffic while maintaining a top speed of 25 km/hr. Mcity in University of Michigan was introduced in 2018 as the first driverless shuttle testing field in the United States. One goal of Mcity project was to analyze how passengers, pedestrians, bicyclists, and other drivers interact with the shuttle and gauge consumer acceptance of the technology [12]. Another autonomous bus project in Oslo was introduced called “smarter transport in the Oslo region (STOR)” [13, 14]. However, the buses were not fully automated they could not distinguish between different objects. In the campus of Aalto University, the driverless shuttle project had operated among other traffic participants for 29 days and 365 km with 522 passengers [15]. Nevertheless, the buses were sensitive to environment dynamics. Snowflakes, heavy rain, and even flying leaves could cause emergency stops. In a comparable endeavor in Finland, two shuttles were provided in this project to cover a 1.5 km trajectory in a park with five stops, interacting with pedestrians and cyclists [16]. Passengers are interviewed to gather their perspectives regarding the utilization of autonomous shuttles [17]. Taiwanese government also examined the potential of autonomous transportation by employing 9-meter driverless buses. These buses were deployed on a designated 2.9 km route in Taichung, Taiwan’s second-largest city, where they successfully tackled tasks such as navigating an open intersection and executing multiple U-turns [18]. UNICARagil was another university-driven project supported by specialists from different enterprises [19, 20, 21, 22, 23, 24]. They created an entirely new vehicle from the ground up and incorporate various enhancements to enable autonomous driving capabilities.
To this end, the contributions in this paper include:
  • A modular software architecture for an autonomous shuttle bus considering adverse weather conditions (rain, snow, fog) which is experimentally validated in real-world scenarios.
  • A perception module with multi-modal sensor fusion for accurate object and snow covered curb detection in adverse weather.
@@
  • A decision-making module, supported by a dedicated safety module, incorporates an intelligent bus stop pullover/merging function specifically tailored for shuttle bus service.
The remainder of this paper is structured as follows. In Section II, WATonoBus sensor selection and calibration are introduced. Section III then presents the WATonoBus software system in details, following by serval case studies and learning discussion in Section IV. Finally, Section V concludes the paper.
@@ II WATonoBus Sensors and Calibration
WATonoBus is equipped with sensors in a configuration as shown in Fig. 1 so that it can sense all surrounding objects. The following subsections will present the work in details.
II-A Sensors and Configuration
The WATonoBus is instrumented with several 3.2MP Basler Ace cameras, a combination of 32 channel Robosense LiDARs, an Applanix POS LVX GNSS, and 2 Continental ARS408-21 Radars to adequately detect any obstacles in its vicinity. The sensor dimensions, mounting positions, reference frames, and transforms are shown in Fig. 2.
@@ II-A Sensors and Configuration
Moreover, the compute system consists of an NVIDIA Jetson Orin AGX embedded unit along with visualization and utility modules on an Intel NUC Ruby with information flow via the Robot Operating System (ROS) [25].
II-B Sensor Calibration
To perform LiDAR-camera calibration, To calibrate LiDAR and camera, we establish a six-degree-of-freedom extrinsic transform. Initially, LiDAR points are projected onto the image plane with a preliminary guess. Refinement is achieved through manual or algorithmic adjustments using projection error, identifiable qualitatively in the image or quantitatively via checkerboard corner positions. This method eliminates the need for specialized calibration boards. Calibration was successfully performed using a developed tool, available here. [^1] https://github.com/Neel1302/LiDAR-camera-calibration.
Figure 2: Illustration of position of the sensors (red) relative to the vehicle body and important distances (in meters). Heights above ground (green) are measured with respect to the road surface. Transformations between sensors are shown in blue.
@@ II-B Sensor Calibration
Calibrating multiple LiDARs is crucial for maintaining a consistent and accurate representation of the environment, aligning data from different LiDARs despite variations in their installation positions and orientations. The Iterative Closest Point (ICP) algorithm is commonly employed for this task [26]. ICP iteratively minimizes differences between point clouds by identifying corresponding points, estimating a transformation matrix for optimal alignment, and repeating this process until a predefined criterion is met. The resulting transformation matrix provides the relative pose between the LiDARs, allowing for accurate 3D perception by bringing all point cloud data into a common coordinate system during operation.
Compensating for ego vehicle motion in LiDAR data processing is crucial to address distortions introduced by the vehicle’s movement during the LiDAR sampling period. For instance, in a system with three LiDARs operating at 10Hz, where each sample takes about 100ms, ego motion can lead to inaccuracies. To mitigate this, a motion compensation method is proposed, leveraging high-frequency Inertial Navigation System (INS) data. INS provides continuous position and orientation information at a higher frequency than LiDARs, enabling real-time correction of LiDAR data. Each point in the point cloud is timestamped, allowing the determination of the LiDAR sensor’s precise pose at the moment of each point’s sampling using INS data. Motion compensation involves establishing the transformation between the recorded pose and the pose at the end of the sampling period. This transformation, applied within 5ms, corrects for ego motion, enhancing the accuracy of the point cloud representation and improving subsequent data analysis and processing tasks.
@@ III WATonoBus Software System
In this section, we propose a multi-modal and modular software architecture with considerations for adverse weather conditions. The high-level system diagram is shown in 3. The following subsections will present each functional module in details.
@@ III-A Perception
Equipped with sensors like cameras, LiDARs, and radars, the bus is adept at navigating in diverse weather conditions. These sensors ensure robust sensing capabilities for the perception module, even in challenging scenarios such as snowy weather.
@@
YOLO (You Only Look Once) v4 stands out as a real-time object detector, showcasing a notable advancement in the YOLO detector series by achieving enhanced accuracy and speed compared to its predecessors [27]. In addition, YOLOv4 adheres to the end-to-end object detection philosophy. This means the model is trained to predict class probabilities and bounding box coordinates directly from full images within one evaluation, distinguishing it from methods involving separate steps for region proposal and classification. The architecture of YOLOv4 compromises three key components: the backbone (CSPDarknet53), the neck (PANet and SAM block), and the head (YOLOv3 head). These components efficiently extract and fuse features, contributing to the model’s performance. Furthermore, YOLOv4 employs various state-of-the-art techniques such as Mish activation, CIOU loss, and DropBlock regularization, alongside features such as multiple anchor boxes per grid cell and three scales for detection. Recognizing these attributes, we developed a ROS package based on YOVOv4 and retrained the network with domain-specific data for our campus environment.
@@ III-A2 LiDAR-only Detection
LiDAR-only detection provides a reliable and accurate enhancement to camera-based detection, particularly for uncommon object and curb detection under adverse weather conditions. It begins with the pointcloud concatenation from the front and two side blind-spot LiDAR sensors. Firstly, an adaptive grid ground segmentation algorithm is employed to remove ground points. Leveraging HD map information, it divides the combined point cloud into multiple regions, allowing for a more accurate and robust estimation of road surfaces. This is achieved by adjusting road surface fitting thresholds based on distances to the ego vehicle, employing stricter thresholds for closer areas and more lenient thresholds for regions further away. This approach effectively enhances the model’s robustness against noises generated by ego vehicle motion or erroneous point cloud reflections, ensuring the detection of all points above the road surface for subsequent object detection. After ground segmentation, an adaptive noise point detection filters out potential points from raindrops or snowflakes. Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm is employed for object clustering [28], enhancing LiDAR-based detection’s robustness in identifying unfamiliar objects. One example for goose detection is shown in Fig. 4. The algorithm identifies obstacle points, selects candidate points close to the ground, and incorporates additional data points in the absence of LiDAR information. For varying curb shapes, a lateral residual model is used. For safety, candidate points on the left of the fitted curb model are reassigned as obstacles, aiding in detecting objects like geese near the curb.
@@
Figure 4: An example for goose detection. Points from the geese are correctly segmented as shown in red, and their corresponding convexhulls are shown in green polylines.
@@
In addition to uncommon object detection, LiDAR-only detection can perceive snow-covered curbs by employing HD maps to partition potential curb points in the point cloud into adjacent road regions. This segmentation provides accurate driving boundary constraints even when actual boundaries and HD maps misalign due to positioning errors or snow obscuration. In challenging scenarios, where snow invades road regions and blurs road boundaries, a weight is introduced during road surface model fitting to mitigate errors. Fig. 5 exemplifies our approach’s effectiveness in curb detection after heavy snow, showcasing a comparison between camera view and point cloud, and revealing actual drivable space detection result in challenging conditions.
@@
Figure 5: (Upper left) Camera view. (Lower left) Point cloud view. Red points means closer distance to ego. (Right) Road boundary detection result after heavy snow. Red points are detected curb points from LiDAR point cloud. Blue squares are estimated road boundary. Yellow line is curb position in HD map. The blue boxes are on the left of the yellow lines, showing the actual drivable space.
@@ III-A3 Radar-only Detection
Radar-only detection offer another level of robustness in WATonoBus perception module since it can operate efficiently under various weather conditions, especially in poor visibility conditions. In addition, the wide filed of view and long-range capabilities (up to 250m) of the selected radar sensors enable anticipatory driving, allowing autonomous vehicles to adjust speed or change lanes in response to upcoming traffic situations. Besides, radar also possesses inherent advantages to distinguish between stationary and moving objects. A customized tracking algorithm is implemented for the selected radar sensors to estimate both longitudinal and lateral velocities of the detected objects. Specifically, longitudinal velocity is essentially the rate of change of distance in the direction of the radar’s line of sight, while lateral velocity refers to the rate of change of distance perpendicular to the radar’s line of sight. Moreover, the objects’ paths can be predicted correspondingly, thus allowing for proactive responses from WATonoBus. These strengths of radar-only detection can complements other sensing modalities, facilitating reliable navigation and safety.
In addition to tracking and prediction, the Radar Cross Section (RCS) of radar-detected objects can be useful in object classification. RCS is a measure of how detectable an object is with a radar. It’s influenced by factors like the size, material, and shape of the object, as well as the radar’s incidence angle. An object with a larger RCS is likely to be a larger vehicle, like a truck or a bus, while an object with smaller RCS might be a motorcycle or a pedestrian. This information can be useful for classifying different objects, thereby enabling ego vehicle to make safer decisions.
Figure 6: A schematic of the overall late fusion algorithm. From the radar firmware, the results are objects with longitudinal and lateral positions and velocities $x_{r},y_{r},v_{x}$, and $v_{y}$. After ground removal and clustering, we have object positions and the dimensions from LiDAR-only detection algorithm, $x_{Li},y_{Li},z_{Li},w,h$, and $l$. From YOLOv4 and camera-based object detection we have bounding box dimensions and class $x_{i},y_{i},w_{i},h_{i}$, $c$. The fusion will complete with the frustum approach and the final output is a bounding box position, dimension, velocity and class $x,y,z,w,h,l,v_{x},v_{y}$.
@@ III-A4 Fusion
This late fusion approach allows us to fully leverage the strengths of each sensor, resulting in a more accurate and reliable object detection and tracking system. The radar fusion with LiDAR point cloud in the first step will enable the scheme to eliminate the false alarms both for radar and LiDAR in challenging weather conditions. Therefore, if the LiDAR-radar fused detections did not find a association with a camera bounding box, it will be labeled as an unknown object. This result combines the best of all sensor capabilities and offers a detailed understanding of each object’s position, dimensions, and motion, enabling the shuttle to make well-informed navigation decisions, thereby significantly enhancing the safety and reliability of WATonoBus.
@@ III-B Localization
WATonoBus requires precise positioning to safely navigate through the drivable space in different weather conditions. This involves accessing the vehicle’s current location and heading while taking into account the location of road boundaries and nearby landmarks through High Definition (HD) maps. Accordingly, a dependable localization module is designed with both GNSS-based and GNSS-denied capabilities.
@@ III-B3 HD Map
To form the HD map, the entire drivable space is discretized into a finite number of equally spaced intervals along an $s$-curve, which is a one-dimensional path coordination system that describes the vehicle’s location with a single scalar (Fig. 7). At each $s$-coordinate, a record of different information is stored in a database that forms the HD map. When WATonoBus is operating in GNSS-denied conditions, the dependable localization module will retrieve the information of reliable landmarks associated with the current s-coordinate from the HD map database.
Figure 7: The HD map contains different layers of information including the relative location of nearby important landmarks (e.g. curb, lane, light poles, building planes, signs, etc.), road type, intersection type, driving condition, etc, stored at each s-coordinate.

### v2

# WATonoBus: Field-Tested All-Weather Autonomous Shuttle Technology

@@ Abstract
All-weather autonomous vehicle operation poses significant challenges, encompassing modules from perception and decision-making to path planning and control. The complexity arises from the need to address adverse weather conditions such as rain, snow, and fog across the autonomy stack. Conventional model-based single-module approaches often lack holistic integration with upstream or downstream tasks. We tackle this problem by proposing a multi-module and modular system architecture with considerations for adverse weather across the perception level, through features such as snow covered curb detection, to decision-making and safety monitoring. Through daily weekday service on the WATonoBus platform for almost two years, we demonstrate that our proposed approach is capable of addressing adverse weather conditions and provide valuable insights from edge cases observed during operation.
@@ I Introduction
Autonomous shuttle bus technology has been steadily advancing in recent years [1]. The introduction of self-driving shuttle buses that can operate along predefined routes without the need for human drivers will enable a new generation of last-mile transportation solutions. Autonomous shuttle buses have the potential to reduce traffic congestion and improve traffic efficiency. By eliminating the need for drivers to operate the vehicle, human errors can minimized and road safety can be enhanced due to the corresponding reduction in collision frequency [2]. Through collaborations among universities, government bodies, and industry partners, numerous cities have initiated pilot programs to test and deploy autonomous shuttle buses in controlled environments [3, 4, 5]. Of these, most are simply exploratory and are motivated by short-term objectives such as the collection of autonomous driving data, the evaluation of economic benefits, and the opportunity to increase public exposure to autonomous driving technology. Although autonomous shuttle buses provide many benefits in comparison to traditional transportation methods, the task of matching or exceeding the performance of a human driver under dynamic weather and lighting conditions remains a challenge. Related safety standards have been drafted to provide qualitative guidelines for testing the performance of autonomous shuttle bus projects [6], however, autonomous shuttle bus projects in adverse weather are still rarely found, posing challenges on the practicability of this technology in such scenarios.
@@ I-A Related Works
In December 1997, Schiphol Airport in Netherlands unveiled one of the early instances of an ”automated people mover” system, known as the ParkShuttle [7]. This pivotal development laid foundation for testing autonomous driving capabilities within a public environment. Currently, numerous global companies have successfully manufactured pilot self-driving shuttle buses for commercial purposes. These firms include Apollo, EasyMile, Navya, and Olli [8, 9, 10]. The WEpod project was initiated in the Netherlands with the aim of enhancing understanding of autonomous vehicles (AVs) [11] where two driverless shuttles navigated through the Wageningen University campus, interacting with traffic while maintaining a top speed of 25 km/hr. Mcity at University of Michigan was introduced in 2018 as the first driverless shuttle testing field in the United States. One goal of the Mcity project was to analyze how passengers, pedestrians, bicyclists, and other drivers interact with the shuttle and gauge consumer acceptance of the technology [12]. Another autonomous bus project in Oslo was introduced called “smarter transport in the Oslo region (STOR)” [13, 14]. However, the buses were not fully automated and could not distinguish between different objects. At Aalto University, a driverless shuttle project operated among other traffic participants for 29 days and 365 kilometers and with 522 total passengers [15]. Nevertheless, the buses were sensitive to environment dynamics. Snowflakes, heavy rain, and even flying leaves caused frequent emergency stops. In a comparable endeavor in Finland, two shuttles covered a 1.5 kilometer path in a park with five stops, interacting with pedestrians and cyclists [16]. The Taiwanese government also examined the potential of autonomous transportation by employing 9-meter driverless buses. These buses were deployed on a designated 2.9 km route in Taichung where they successfully tackled tasks such as navigating an open intersection and executing multiple U-turns [17]. UNICARagil was another university-driven project supported by specialists from different enterprises [18, 19, 20, 21, 22]. They created an entirely new vehicle from the ground up and incorporated various enhancements however did not address adverse driving scenarios.
To this end, we aim to make the following contributions:
  • A modular software architecture for an autonomous shuttle bus, first of its kind in Canada [23], considering adverse weather conditions (rain, snow, fog) which is experimentally validated in real-world scenarios.
  • A perception module with multi-modal sensor fusion for accurate object and snow-covered curb detection in adverse weather.
@@
  • A decision-making module, supported by a dedicated safety module, incorporating an intelligent bus stop pullover/merging function specifically tailored for shuttle bus service.
The remainder of this paper is structured as follows. In Section II, the WATonoBus sensor suite is introduced. Section III then presents the WATonoBus software system, followed by several case studies and discussion on insights in Section IV. Finally, Section V concludes the paper.
@@ II WATonoBus Sensors and Calibration
WATonoBus (video) is equipped with sensors placed in a configuration that allows sensing all surrounding objects as shown in Fig. 1. WATonoBus is instrumented with 6 3.2MP Basler Dart cameras, three 32 channel Robosense LiDARs, an Applanix POS LVX GNSS, and two Continental ARS408-21 radars to adequately detect any obstacle in its vicinity.
@@
Moreover, the compute system consists of a NVIDIA Jetson Orin AGX embedded unit along with visualization and utility modules on an Intel NUC Ruby with information flow via the Robot Operating System (ROS) [24].
@@ III WATonoBus Software System
In this section, we propose a multi-modal and modular software architecture with considerations for adverse weather conditions. A high-level system diagram is shown in Fig. 2.
@@ III-A Perception
You only look once (YOLO) v4 stands out as a real-time object detector, showcasing a notable advancement in the YOLO detector series by achieving enhanced accuracy and speed compared to its predecessors [25]. The architecture of YOLOv4 comprises of three key components: the backbone (CSPDarknet53), the neck (PANet and SAM block), and the head (YOLOv3 head). These components efficiently extract and fuse features, contributing to the model’s performance. Furthermore, YOLOv4 employs various state-of-the-art techniques such as Mish activation, CIOU loss, and DropBlock regularization, alongside features such as multiple anchor boxes per grid cell and three scales for detection. Recognizing these attributes, we developed a YOVOv4 ROS package and retrained the network with domain-specific data for our campus environment (code).
@@ III-A2 LiDAR-only Detection
LiDAR-only detection provides a reliable and accurate enhancement to camera-based detection, particularly for uncommon object classes and curb detection under adverse weather conditions. It begins with the point cloud concatenation of the front and two side blind-spot LiDARs. Firstly, an adaptive grid ground segmentation algorithm is employed to remove ground points. Leveraging high definition (HD) map information, we divide the combined point cloud into multiple regions, allowing for a more accurate and robust estimation of road surfaces. This is achieved by adjusting road surface fitting thresholds based on distance from the ego vehicle, employing stricter thresholds for closer areas and more lenient thresholds for regions further away. This approach effectively enhances the model’s robustness against noise attributed to ego vehicle motion or erroneous point cloud reflections, ensuring the detection of all points above the road surface for subsequent object detection. Post ground segmentation, an adaptive noise point detection algorithm filters out raindrops or snowflakes (video 1 and 2). Density-based spatial clustering of applications with noise (DBSCAN) algorithm is employed for object clustering [26], enhancing LiDAR-based detection’s robustness in identifying unfamiliar objects. An example for geese detection is shown in Fig. 3 (video 1 and 2). The algorithm identifies obstacle points, selects candidate points close to the ground, and incorporates additional data points in the absence of LiDAR information. For varying curb shapes, a lateral residual model is used. For safety, candidate points on the left of the fitted curb model are reassigned as obstacles, aiding in detecting objects like geese near the curb.
@@
Figure 3: An example for geese detection. Points from the geese are correctly segmented as shown in red, and their corresponding convex hulls are shown in green polylines.
@@
In addition to uncommon object detection, LiDAR-only detection can perceive snow-covered curbs by employing HD maps to partition potential curb points in the point cloud into adjacent road regions. This segmentation provides accurate driving boundary constraints despite a mismatch in the HD map information due to positioning errors or due to snow. In scenarios where snow covers the curbs and leads to tighter road boundaries, a weight factor is introduced during road surface model fitting to mitigate errors. Fig. 4 exemplifies our approach’s effectiveness in curb detection in heavy snow, showcasing a comparison between camera view and point cloud, revealing drivable space detection results (video).
@@
Figure 4: (Upper left) Camera view. (Lower left) Point cloud view. Red points denote closer distance to ego. (Right) Road boundary detection results in heavy snow. Red points are detected curb points from LiDAR point cloud. Blue squares are estimated road boundary. Yellow line denotes curb position from HD map. The blue boxes are on the left of the yellow line, reflecting actual drivable space.
@@ III-A3 Radar-only Detection
Radar-only detection offers another level of robustness in the perception module since it offers reliable operation under various weather conditions, especially in poor visibility conditions. In addition, the wide field of view and long-range capabilities (up to 250m) of the selected radar sensors enable anticipatory driving, allowing autonomous vehicles to adjust speed or change lanes in response to upcoming traffic. Radar also possesses inherent advantages to distinguish between stationary and moving objects.
In addition to tracking and prediction, the Radar Cross Section (RCS) of radar-detected objects can be useful in object classification. This information can be useful for classifying different objects, thereby enabling the ego vehicle to make safer decisions.
Figure 5: A schematic of the overall late fusion algorithm. The radar provides longitudinal and lateral positions and velocities $x_{r},y_{r},v_{x}$, and $v_{y}$ of objects. After ground removal and clustering, we have object positions and the dimensions from LiDAR-only detection algorithm, $x_{Li},y_{Li},z_{Li},w,h$, and $l$. From YOLOv4 and camera-based object detection we have bounding box dimensions and class $x_{i},y_{i},w_{i},h_{i}$, $c$. The fusion process completes with frustum association and the final output is bounding box position, dimension, velocity and class $x,y,z,w,h,l,v_{x},v_{y}$ for each object.
@@ III-A4 Fusion
This late fusion approach allows us to fully leverage the strengths of each sensor, resulting in a more accurate and reliable object detection and tracking system. The radar fusion with LiDAR point cloud in the first step will enable the scheme to eliminate the false alarms both for radar and LiDAR in challenging weather conditions. Therefore, if the LiDAR-radar fused detections did not find a association with a camera bounding box, it will be labeled as an unknown object. This result combines the best of all sensor capabilities and offers a detailed understanding of each object’s position, dimensions, and motion, enabling the shuttle to make well-informed navigation decisions, thereby significantly enhancing the safety and reliability of WATonoBus (video).
@@ III-B Localization
WATonoBus requires precise positioning to safely navigate through the drivable space in different weather conditions. This involves accessing the vehicle’s current location and heading while taking into account the location of road boundaries and nearby landmarks through HD maps. Accordingly, a dependable localization module is designed with both GNSS-based and GNSS-denied capabilities.
@@ III-B3 HD Map
To form the HD map, the entire drivable space is discretized into a finite number of equally spaced intervals along an $s$-curve, which is a one-dimensional path coordination system that describes the vehicle’s location with a single scalar. At each $s$-coordinate, a record of different information is stored in a database that forms the HD map. When WATonoBus is operating in GNSS-denied conditions, the dependable localization module will retrieve the information of reliable landmarks associated with the current s-coordinate from the HD map database.